Status and Schedule

Data refreshes automatically every 5 minutes. Click an image to enlarge.

Recent Status Messages

Want to receive APS status notifications automatically?
Aug, 20 2014 11:05:57 Beam expected ~12:30
Due to unforeseen issues with the injector, we are extending the beam estimate another hour to ~12:30.
Aug, 20 2014 10:09:23 Beam Loss at 09:24, recovery ~11:30
Stored beam was lost at 09:24 due to S11 EPS vacuum fault, recovery is expected to take until ~11:30 to finish S11 repairs and recover the injector.
Aug, 18 2014 09:09:39 SR pump switch successful
The SR water pump switch is complete, there remain 100mA stored beam.
Aug, 18 2014 08:32:36 SR pump switch at 09:00
At 09:00, we are performing a water pump switch due to a leak. We hope it will be bumpless, but there is a risk that stored beam will be lost. In the event of a dump, recovery time will be ~45-60minutes.
Aug, 14 2014 22:19:57 Shutter Permit Enabled
At 22:17 shutter permit was enabled with 102mA of stored beam.
Aug, 14 2014 22:05:51 Stored Beam Lost Update
We anticipate on enabling shutters at ~22:25.
Aug, 14 2014 21:28:35 Stored Beam Lost
At 21:24 stored beam was lost. We are investigating.
Aug, 11 2014 19:02:30 Beam Motion and Orbit Correction for Horizontal Plane Suspended from 6:38PM to 6:50PM.
On August 11, 2014, at 6:38PM, we experienced minor beam motion and Orbit Correction for the Horizontal plane was suspended from 6:38PM to 6:50PM. The issue has been resolved and Orbit Correction is running again. -Ops.
Aug, 11 2014 14:25:32 Shutter permit disabled at 14:23
Shutter permit was disabled at 14:23. Stored beam was lost. An investigation is ongoing.
Aug, 11 2014 10:55:49 Shutter permit enabled at 10:52
On Monday, August 11th, Shutter permit was enabled at 10:52. There is ~102mA of stored beam for Fill #33. Top-up is ongoing.
Aug, 11 2014 10:31:31 Shutter permit disabled at 10:22
Shutter permit was disabled on Monday, August 11th at 10:22. Stored beam was lost. Fill #32 is over. First latch fault was S9 Water/Vac. [X]
Aug, 09 2014 12:11:05 Shutters enabled
On August 9, 2014 at 11:56AM, Shutter permit was enabled with 102mA of stored beam, Top-Up is ongoing.
Aug, 09 2014 11:08:53 29/30 difficulties
1st attempt at repair of S29/30 MPS issue was unsuccessful, the engineer on call is en route to assist with further troubleshooting & repairs. Estimate for beam is unknown at this time, updates will be forthcoming.
Aug, 09 2014 10:29:46 29/30 repairs started
Technicians have arrived ~10:20 and are starting repairs of the 29/30BPLD issue. If successful, beam should be back within the hour.
Aug, 09 2014 08:58:15 Beam Loss 08:16, 2hr ETA
Stored beam was lost at 08:16 Sat, 8/9 due to S29/30BPLD trip that will not reset. Personnel are being called in to troubleshoot/repair. Next estimate in ~2hours at 10:16.
Aug, 08 2014 18:05:40 Top-Up Resumed
On August 8, 2014, at 5:43PM, Top-Up was aborted after communications were lost to an IOC. Top-Up was resumed at 6:03PM. -Ops.
Aug, 08 2014 17:46:04 Top-Up aborted due to IOC issue.
On August 8, 2014, at 5:43PM, Top-Up was aborted after communications were lost to an IOC. - Ops
Aug, 08 2014 00:45:21 Top-Up resumed with 102mA stored beam
Shortly before midnight repairs were completed on the failed Linac RF system. The injectors were recovered and Top-Up restarted by midnight. A fill-on-fill was performed at 00:30 and Top-Up is ongoing.
Aug, 07 2014 22:37:21 Top-Up aborted at 20:47 due to Linac Injector issue.
On August 7, 2014 Top-Up was aborted at 20:47 due to a Linac injector issue. RF Group technicians have arrived to the APS and are starting repairs. As of now, there is no estimated time for Top-Up. A Fill-On-Fill will be scheduled when repairs are complete. Next update in ~ 1 hour. -Ops.
Aug, 07 2014 21:59:24 Top-Up aborted at 20:47 due to Linac Injector issue.
On August 7, 2014 Top-Up was aborted at 20:47 due to a Linac injector issue. RF Group technicians have been called in for repairs and are en route to the APS. As of now, there is no estimated time for Top-Up. A Fill-On-Fill will be scheduled when repairs are complete. Next update in ~ 1 hour. -Ops.
Aug, 07 2014 20:54:52 Topup aborted at 20:47 due to Linac injector issue
Topup was aborted at 20:47 due to Linac injector issue. No estimated time for topup at this time
Aug, 07 2014 17:19:42 Shutters Enabled
On August 7, 2014 Stored beam was lost at 04:07PM. Shutter Permit was enabled at 05:16PM with ~102mA of stored beam. Top-Up is ongoing. -Ops.
Aug, 07 2014 16:18:08 Stored beam lost
On August 7, 2014 Stored beam was lost at 04:07PM, we are investigating, Next Fill Info ETA ~ 5:15PM. -Ops.
Aug, 06 2014 16:50:21 OrbCorr/RTFB restarted 16:49
The S12/13 orbit bump has been fixed and orbit correction with RT feedback was restarted 16:49.
Aug, 06 2014 16:45:53 Orbit corr/RTFB off 16:40
To address a bump in S12/13 post the earlier reboot of 11VP, orbit correction & RT feedback have been aborted.
Aug, 06 2014 16:27:33 Top-up is ongoing 16:20
IOCS11VP reboot/recovery is complete, top-up and orbit correction with RT feedback are running as of 16:20.
Aug, 06 2014 15:43:06 Top-up aborted 15:42
Top-up was aborted at 15:42 in order to reboot iocs11vp, orbit correction and RTFB will also be off until reboot & recovery are complete.
Aug, 01 2014 08:55:18 Top-Up
On August 1, 2014, at 08:48AM , Top-Up was aborted due to an injector issue, at 08:52AM, Top-Up was restarted.